OpenStudy (anonymous):

maybe add more symbols that have to do with math like the "sines" and what not or even openstudys own little dictionary of terms

OpenStudy (♪chibiterasu):

\[\sin \cos \tan\]

Nnesha (nnesha):

\[\huge\rm \theta \] \theta\[\large\rm 56^o\] ^o degree